# On the Formalization of Cryptographic Migration

### v3

@@ Abstract
With the advancement of quantum computing, the transition to post-quantum cryptography (PQC) is becoming increasingly critical to maintain the security of modern dependable infrastructural systems. This paper presents a novel approach to gain insight into the structure of cryptographic migration problems, using a semi-formal model to capture the inherent dependencies and complexities of such transitions. Using classical mathematical results from combinatorics, probability theory, and combinatorial analysis, we assess the challenges of migrating large cryptographic IT-infrastructures and prove that – in a suitable sense – cryptographic migration has a certain expected complexity. Furthermore, we analyze the proposed model in terms of real-world patterns as well as its practical applicability, and discuss difficulties that arise when trying to model real-world migration projects. This work sets the stage for future advances in both the theoretical understanding and practical implementation of cryptographic migration strategies in the post-quantum era.
@@
Cryptographic migration, post-quantum cryptography, semi-formal model
@@ I Introduction
Cryptography is fundamental to ensuring the information security in today’s digital world and is thus one important cornerstone of many modern dependable systems. Despite advances in cryptographic technologies, the process of migrating existing systems and IT-infrastructures to new standards presents significant challenges [3]. One example is the transition from IPv4 to IPv6, which took more than 20 years and promoted technologies such as gateways and NAT to circumvent the actual migration instead. A cryptographic example is the update of key sizes (RSA keys got bigger and so did the integration problems) or the transition between different hash functions (SHA-1 to the SHA-2 family to the SHA-3 family) with slow adoption of the SHA-3 family as SHA-2 seems still sufficient for most tasks. It seems that a rapid transition requires a strong incentive in terms of the benefits of adopting the new technology. One notable example of quicker migration, at least in the broad public, is the transition from the Data Encryption Standard (DES) to the Advanced Encryption Standard (AES), which was driven by the need for stronger security measures due to DES’s vulnerability to brute force attacks [20].
The issue of migration is particularly prominent in the current transition towards PQC, which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the work-hypothesis for the high-security sector that there likely exists a cryptographically relevant quantum computer in the early 2030s [8]. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g. classified) data often has long confidentiality requirements (e. g. 25 years). In addition, migrating the IT-infrastructure correspondingly takes a considerable amount of time [24]. This runs against estimates on how long it will take to build a crypto analytically relevant quantum-computer, c. f. Mosca’s famous inequality [13].
@@
To be able to conduct such migrations as efficiently as possible we aim to model the PQC migration problem mathematically and look for potential optimal strategies. As two relevant parts, i.e. the modeling of IT networks and dependencies, are often depicted as graphs [11], we also examine the migration problem in the context of a graph-based model. By this systematic analysis, we hope to better understand the structural problems and find strategies necessary for building tools to assist with conducting the actual migrations. Our findings show that usually the systems’ dependencies of a to-be-migrated infrastructure necessitate a compatible migration order, otherwise functionality is lost. Within this order there are always sets of systems which need to be migrated together – so called migration clusters. As soon as the systems under evaluation exhibit non-trivial complexity, we have many or large clusters, and these clusters are expected to be interdependent, making migration difficult. To facilitate the process and still allow for rapid migration, we see a need for tools that build on our results to find and visualise the optimal migration graphs.
After an overview of related work in Section II, we present our first contribution in Section III. It consists of a formalization approach for cryptographic migration, a discussion of the architectural structure of the resulting model, and explores typical patterns that occur in practice. In Section IV we prove that any (reasonably large) migration project has an expected intrinsic complexity as our second contribution. Section V further describes our model in terms of its usage and its potential applications which shall serve as a foundation for tools which assist the upcoming real-world migrations – our third contribution.
@@ II Background and Related Work
After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [1], the transition to PQC has started in practice: in the US, there is a project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [3]. In Europe, the German BSI published a guideline concerning post-quantum migration [7] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [9]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [2].
@@
Despite these world-wide activities, we found little research on both, the general structure of migration projects and how to suitably model cryptographic migration. In fact, many people consider migrating of soft- and hardware components as a best-practice task. So, people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [18]). Shouldn’t there be structural reasons why cryptographic migration is so hard to achieve in practice? We will show in Section IV that there is!
Recent efforts have focused on consolidating the state-of-the-art in cryptographic migration [14] and systematically organizing knowledge on cryptographic agility [16]. Furthermore, frameworks arise that aim to support the migration process as in [25], and the development of tools to assist in the migration process [11]. The latter follows – as we do – a graph-based approach, providing a solid starting point for PQC migration by introducing a practical framework focused on cryptographic and data inventories and dependency analysis within enterprise systems. However, the analysis in [11] primarily addresses operational aspects and could benefit from a more formal approach. Building on this groundwork, our efforts enhance the theoretical rigor by formalizing migration challenges through mathematical models. Generalizing the approach might even address broader migration scenarios beyond PQC-specific contexts.
Also to be mentioned are few informal models of the migration problem (e. g. [19, 15, 23, 6]). However, formal treatment is, as far as the authors know, still pending. We believe that our migration model together with its analysis can substantially contribute to the formal basics of cryptographic migration.
@@ III A Model for the Cryptographic Migration Problem
Any IT-infrastructure we migrate consists of a number of migration components or simply components which need to be migrated. Such migration components can be software libraries, whole software packages, full operating systems, network protocols, etc. just to name a few. For a more thorough study of which components we might encounter in practice see Section III-B. The migration problem is now to migrate all migration components. However, not every component can be migrated at every point in time, because there might be dependencies between the components. For example, a given software package can only be migrated to post-quantum if its used cryptographic libraries can. Another example is the migration of a communicating software component, which can only be migrated if the corresponding counterparts of the communication are migrated at the same time. In this case the migration can be done (if there are no other dependencies) if all communicating parties of this single migration component are migrated at the same time. We call this as stated in the introduction in Section I a migration cluster. Components without dependencies outside its migration cluster are called migratable components. The migration process looks now as follows:
@@
    In every migration step we select, following a suitable migration strategy, a migratable component.
@@
    We then compute its migration cluster, i.e. possible interdependent components that need to be migrated at the same point in time.
@@
    This process is repeated until there is no migration component left and we have then solved the migration problem for this specific IT-infrastructure.
The whole model is now captured in a directed graph whose vertices are the migration components and the edges are the dependencies between components, the migration graph of the migration problem, see Appendix -A for the formal details.
@@
Note that a migration graph’s structure will vary significantly depending on the concrete IT-infrastructure and the level of model we have in mind, c.f. Section III-A. But also the abstraction layer and the type of migration lead to many design choices of the model. The resulting graphs will show parallels to those that already exist in several disciplines of computer science. Those comprise flowcharts, call graphs, and finite-state machines in programming, network diagrams, or software dependency graphs, just to name a few. Additionally, current efforts in the creation of Software Bills of Materials (SBOMs) or Cryptography Bills of Materials (CBOMs) and their respective tooling add to that collection of suitable graphs.
We expect that post-quantum migration is the most significant type of cryptographic migration of IT-infrastructures in the near future [14]. Its specific properties introduce aspects that were less prevalent in the classical cryptography world. One example are hybrid cryptographic schemes. Depending on the specific model, a hybrid solution may just be seen as a single cryptographic primitive or cipher suite. Yet, modeling detailed software dependencies or taking dependencies on space/throughput requirements into account might have a significant effect.
@@
Post-quantum migration also emphasizes the importance of cryptographic agility (see for example [18]). Many communication protocols used to be optimized for a limited set of somewhat similar algorithms. The introduction of crypto-agile mechanisms that support various kinds of cryptographic primitives and, sometimes arbitrary, combinations of these, will lead to an increase in complexity. In a graph which only models systems and their protocol-related communication relations in between, one may only treat this implicitly. Yet again, while graphs that take other dependencies into account usually become harder to model, they sometimes better reflect the complicated circumstances of the real world.
Also, regulatory requirements of post-quantum solutions and different views on current and future algorithm recommendations may become significantly more complicated, especially in a multinational context.

### v4

# rReviewed Literature

@@ Abstract
We present a novel approach to gaining insight into the structure of cryptographic migration problems which are classic problems in applied cryptography. We use a formal model to capture the inherent dependencies and complexities of such transitions. Using classical mathematical results from combinatorics, probability theory, and combinatorial analysis, we evaluate the challenges of migrating large cryptographic IT infrastructures and prove that – in a suitable sense – cryptographic migration exhibits a certain expected complexity. We also provide numerical data for selected parameter sets. Furthermore, we analyze the proposed model in terms of real-world patterns and its practical applicability. Additionally, we discuss the challenges of modeling real-world migration projects. As concrete examples we examine the transition to post-quantum cryptography of the CI/CD system GitLab and the multi-level technological transition of distribution power grids. This work paves the way for future advancements in both the theoretical understanding and practical implementation of cryptographic migration strategies.
@@
cryptographic migration, formal model, hardness result, post-quantum transition
@@ I Introduction
Cryptography is fundamental to ensuring information security in today’s digital world and is thus an essential cornerstone of many modern dependable systems. Despite advances in cryptographic technologies, the process of migrating existing systems and IT infrastructures to new standards presents significant challenges [1]. One example is the transition from IPv4 to IPv6, which took more than 20 years and promoted technologies such as gateways and network address translation (NAT) to circumvent the actual migration. In the cryptographic realm, notable examples include updating key sizes (where RSA key sizes increased, as did integration problems) and transitioning between different hash functions (from SHA-1 to the SHA-2 family, and eventually to the SHA-3 family), with the SHA-3 family being adopted slowly as SHA-2 is still considered sufficient for most tasks. Rapid transition appears to require a strong incentive to adopt new technology. A notable exception is the transition from the symmetric cipher DES to its replacements AES, driven by the need for stronger security measures due to DES’s vulnerability to brute-force attacks [2]. The issue of migration is particularly prominent in the current transition towards post-quantum cryptography (PQC), which aims to protect data against future quantum computing threats. The German Federal Office for Information Security (BSI) uses the working hypothesis for the high-security sector that there likely exists a cryptographically relevant quantum computer in the early 2030s [3]. Mosca’s famous inequality [4] states that there is a 1 in 2 chance that such a computer will exist in 2031. Thus, there are numerous practical post-quantum migrations going on world-wide. The reason is that we are already running-out of time: Sensitive (e. g., classified) data often has long confidentiality requirements (e. g., 25 years) and it might be possible to decrypt cryptographically protected data which is recorded today in the future. In addition, migrating the IT-infrastructure correspondingly takes a considerable amount of time [5].
Our approach to support PQC migration is to model the migration problem suitably and identify this way optimal migration strategies in general. We propose a systematic analysis to better understand the structural problems and explore which potential models are necessary for building tools to assist actual PQC migrations. Although algorithmic aspects and the details of the concrete tooling are beyond the scope of this work, we believe that a structured assessment of the migration problem provides valuable insights into the structural complexties involved.
@@
Our findings show that the dependencies of a system to be migrated usually necessitate a compatible migration order to avoid losing functionality. Within this order, there are always sets of systems which need to be migrated together – so called migration clusters (cf. Definition 1) – and their successive migration defines the corresponding migration strategy. As soon as the systems under evaluation exhibit non-trivial complexity, we have many or large clusters, and these clusters are expected to be interdependent, making parallelization of the different steps impossible and migration thus difficult. To facilitate the process while still allowing for rapid migration, we need tools that can identify and visualize optimal migration strategies. In order to cope with the complexity of developing such tools, we believe that a proper understanding of the underlying problem is necessary. Our work presented here provides the formal basis for this.
Contributions
Our contributions are as follows:
  1. 1.
    We introduce a novel graph-based formalization of cryptographic migration, along with a discussion of the structure of the resulting model (Section III).
  2. 2.
    We then prove that any (reasonably large) migration project has an expected complex dependency structure and thus show that this structure prohibits parallelizing the migration (Section IV).
  3. 3.
    We explore typical graphs and patterns that occur in practice and how these patterns can support system administrators in planning a migration project (Section V).
  4. 4.
    We demonstrate how our model can be applied to an actual migration problem as an exemplar to showcase its ability to draw conclusions about real-world migrations. We also discuss how multiple graphs on different levels of abstraction can be combined when planning the migration of a complex critical infrastructure system (Section VI).
Overall, the formalization proposed in this paper is an important step toward a better, more systematic understanding of migration problems. It also lays the groundwork for developing tools that will support engineers in the complex cryptographic migrations that lie ahead in the coming years.
@@ II-A Background
After the announcements of the first post-quantum public key algorithms for consideration as an international standard by the National Institute of Standards and Technology (NIST) [6, 7, 8], the transition to PQC has started in practice: in the US, there is a project run by NIST with a large consortium of industry-leading companies that explores the transition to PQC [1]. In Europe, the German Federal Office for Information Security (BSI) published a guideline concerning post-quantum migration [9] and also included certain post-quantum algorithms in their recommended suite of cryptographic algorithms [10]. The Dutch government also published a “PQC Migration Handbook” containing guidelines for migrating to PQC [11].
@@
Despite these world-wide activities, we found little research on the general structure of migration projects and how to suitably model cryptographic migration. In fact, many people consider migrating soft- and hardware components as a best-practice task. So, people legitimately ask “Where Is the Research on Cryptographic Transition and Agility?” as David Ott did at RWC2022 (see also [12]). We show in Section IV that there are structural reasons why cryptographic migration is so hard to achieve in practice.
Ott et al. define the term cryptographic agility as “the ability to make a transition from one cryptographic algorithm or implementation to another” [12]. Our own, canonical definition [13] states that “Cryptographic Agility is a theoretical or practical approach, objective, or property which provides capabilities for setting up, identifying, and modifying encryption methods and keying material in a flexible and efficient way while preserving business continuity.” Adopting cryptographic agility in this sense ensures the flexibility to transition between algorithms or implementations in response to evolving threats, see Section V-B for more details. Thus, in order to implement cryptographic agility we need to prepare cryptographic migration. In turn, for preparing migration, we need a step-by-step migration plan which must allow careful re-evaluation of dependencies and system integrity at each stage. This kind of approach is in line with making security solutions sustainable by allowing to adapt to changing threats and technologies.
II-B Related Work
Our recent efforts have focused on consolidating the state-of-the-art in cryptographic migration [14] and systematically organizing knowledge on cryptographic agility [13]. Furthermore, frameworks arose that aim to support the migration process as in [15]. Also, there are few informal models of the migration problem (e. g., [16, 17, 18, 19]). All these models have in common that they address operational aspects of cryptographic migration in various institutional contexts. This can be aligned with the above observation of structuring best-practice tasks. While these kind of analyses are clearly important and can serve as a basis for practical migration, they do not attempt to formally assess the underlying structural problem. Instead they give guidelines on how to assess the needed transition efforts, explain how to perform a suitable risk-analysis, and describe process landscapes to perform cryptographic transition.
One way of collecting the information that is necessary to build a migration model is to construct a software bill of materials (SBOM) [20] or a cryptographic bill of materials (CBOM) [21, 22]. While the SBOM contains information about which software in which version is used and can also include information about dependencies that are relevant for cryptography, the CBOM contains information about the concrete cryptographic mechanisms in use. This information can be used in conjunction to plan a migration, e. g., by identifying vulnerable protocols in the communication between two assets in the CBOM and then using the SBOMs of these assets to locate the software that should be migrated.
Building an SBOM or a CBOM requires extensive tooling for (semi-)automated collection. There is a plethora of relevant tools available111see for example https://perma.cc/N8WE-QBTZ that help either on the system level or on the network level to collect the desired information of the infrastructure. The results are typically vast: even for small setups, such an automated assembly can easily yield tens of thousands of components.
However, complex migration projects require many successive steps and in-depth strategic planning. Based on an SBOM/CBOM alone, only a simple risk-assessment and migration of components – without formal structure – based only on the highest risk exposure or when updates are available is feasible. We believe this is insufficient since it ignores the complexity of migration projects that should be based on more information than the SBOM and CBOM provide, including information about exposure, and should be based on a structured plan with a comprehensive overview rather than on ad-hoc interventions.
Initial standardization attempts with the same aim include frameworks like CARAF, a risk management approach for post-quantum migration [23]. In this framework, risks are identified, an asset inventory is set up and the risk is estimated for each asset, yielding an organizational road-map. Relating that to our migration model, after the asset inventorization we also get a list of relevant components (i. e., the nodes within the graph), but the CARAF approach does not provide any information about the dependencies between the components.
The draft “Repeatable Framework for Quantum-safe Migrations” of the framework ETSI published [19] approaches the migration problem from an organizational perspective. It assumes that the responsibility for migration projects is assigned internally while some form of governance regulates the interdependence of the parties involved. Local dependency structures are described in part, but an analysis of longer chains of dependencies is not conducted. Therefore, only local dependencies can be captured by this framework.
As far as we know, a formal treatment for the migration problem is still pending: While Hassim et al. [24] introduce a solid starting point for a PQC migration model by analyzing certain local dependencies in real-world migration projects, these authors do not work out the resulting composed dependency structures found in real-world interconnected systems. Instead they provide a practical framework focused on cryptographic and data inventories and a local dependency analysis within enterprise systems. Also, their analysis primarily addresses operational aspects.
Building on this groundwork, our efforts enhance the theoretical rigor by formalizing migration challenges through mathematical models. Our generalization even addresses broader migration scenarios beyond PQC-specific contexts. We believe that our migration model together with its analysis and study of real-world applicability can substantially contribute to the understanding of the formal foundations of cryptographic migration.
In any case, combining different kinds of information, carving out relations and creating a model of the migration graph is still a lot of manual work and is currently not fully supported by tooling. In future work we envision tooling to support the creation and maintenance of migration graphs as well as their analysis.
@@ III A Model for the Cryptographic Migration Problem
Having motivated the need for a rigorous treatment of cryptographic migrations, we formulate a migration project as a directed graph whose vertices represent concrete software, hardware or organizational components which need to be migrated. Its edges capture dependencies that constrain the order of those migrations. We introduce an intuitive model, moving to a precise definition that introduces the key notions of strongly-connected migration clusters, migratable components, and the scheduling strategies that remove such clusters step-by-step. This formal framework sets the stage for the complexity analysis in Section IV and underlies the practical patterns developed later in the paper.
III-A Informal Description of the Model
Any IT infrastructure we migrate consists of a number of migration components or simply components which need to be migrated. For an overview of which components we might encounter in practice see Section V-C. The migration problem is now to migrate all migration components. However, not every component can be migrated at every point in time, because there might be dependencies between the components. For instance, an application can migrate to post-quantum algorithms only after the cryptographic libraries it relies on have been upgraded. Another example is the migration of a communicating software component, which can only be migrated if the corresponding counterparts of the communication are migrated at the same time or if we introduce some kind of interoperability. In this case the migration can be done (if there are no other dependencies) if all communicating parties of this single migration component are migrated at the same time. We call this a migration cluster. Components without dependencies outside its migration cluster are called migratable components. The migration process looks as follows:
@@ III-A Informal Description of the Model
    In every migration step we select a migratable component, following a suitable migration strategy.
@@
    We then compute its migration cluster, i. e., possible interdependent components that need to be migrated at the same point in time.
@@
    This process is repeated until there is no migration component left and we have then solved the migration problem for this specific IT infrastructure.
@@
The whole model is now captured in a directed graph whose vertices are the migration components and the edges are the dependencies between components, the migration graph of the migration problem, which we will detail in this section. Note that a migration graph’s structure will vary significantly depending on the concrete IT infrastructure and the kind of model we have in mind. But also the abstraction layer and the type of migration lead to many design choices of the model, cf. Section V. The resulting graphs will show parallels to those that already exist in several disciplines of computer science such as flowcharts, call graphs, and finite-state machines in programming, network diagrams, or software dependency graphs.
@@
We expect that post-quantum migration is the most significant type of cryptographic migration of IT infrastructures in the near future [14]. Its specific properties introduce aspects that were less prevalent in the classical cryptography world, where examples are hybrid cryptographic schemes and more complex instantiations of cryptographically agile solutions. In a graph which only models systems and their protocol-related communication relations in between, one may only treat this implicitly. Yet again, while graphs that take other dependencies into account usually become harder to model, they sometimes better reflect the complicated circumstances of the real world.